Carl Lewis , a sprinter in the 1988 Olympic Games, ran the 100 meter dash in 8.90 sec. what was his speed in miles per hour ?


1
Expert's answer
2021-09-09T00:27:02-0400

Speed=D/T

100÷8.90

=11.236 m/s

1m/s=2.237miles per hour

11.236m/s to miles per hour=11.236×2.237


=25.135 miles per hour